# Modular Machine Learning: An Indispensable Path towards New-Generation Large Language Models

### v1

@@ 3 The Key Significance of MML for LLMs
Enable Counterfactual Reasoning. One of the most profound limitations of current LLMs is their inability to effectively reason about counterfactuals [2]. Counterfactual reasoning involves exploring hypothetical scenarios that differ from the observed reality to answer “what if” questions. For example, in medicine, counterfactual reasoning might explore the outcome which assumes a patient had received a different treatment. While LLMs excel at predicting texts based on patterns in large datasets, their statistical foundations make them ill-suited for tasks requiring logical consistency and causal reasoning. MML bridges this gap by integrating modular reasoning systems capable of counterfactual reasoning. These systems rely on explicitly defined rules or causal graphs to model relationships between variables, enabling the generation of logically coherent hypotheses about unobserved scenarios. By non-trivially integrating such modules into LLMs, it is possible to enhance the ability of LLMs to handle tasks requiring robust logical reasoning [3]. The integration can take various forms. One approach is to use neural modules as perception engines to parse and preprocess complex, high-dimensional data into structured representations compatible with modular systems. These modular systems then perform counterfactual reasoning by manipulating these structured representations according to well-defined logical or causal principles. For example, in policy decision-making, an MML-enhanced LLM could analyze historical data to suggest alternative outcomes under different policy scenarios, offering interpretable insights for potential causal mechanisms. Beyond applications, MML-based frameworks can provide architectural advantages by separating perception from reasoning. This modular design ensures that the reasoning layer operates independently of the statistical biases inherent in neural networks. For instance, while a neural network might overfit correlations from training data, an MML-based module is able to enforce constraints that align with known causal relationships, mitigating the risk of generating spurious counterfactuals. Thus, MML could serve as a foundation for counterfactual reasoning within LLMs [4].
@@
Eliminate Hallucinations. A critical challenge for LLMs is the phenomenon of hallucinations—the generation of plausible yet factually incorrect or inconsistent information [6]. Hallucinations stem from the probabilistic nature of LLMs, which are trained to optimize likelihood functions rather than ensure factual correctness. This limitation poses significant risks in high-stakes domains such as healthcare, legal systems, scientific research, etc. MML offers a transformative solution by introducing modular learning and reasoning modules that enforce factuality. These modules operate as logical evaluators, cross-referencing LLM outputs with structured knowledge bases, ontologies, or predefined rules. For instance, in a medical diagnosis application, modular systems could validate treatment recommendations generated by an LLM against established clinical guidelines, flagging any inconsistencies. Furthermore, MML can enable iterative refinement of LLM outputs through modular post-processing. Consider a scenario where an LLM generates a scientific summary. An MML-based module could validate the summary by applying logical rules derived from domain-specific knowledge, such as consistency checks for units, measurements, or causal relationships. This iterative process ensures that the final output is both factually accurate and logically coherent. Another promising direction is the use of MML for real-time monitoring and correction of LLM outputs. MML-based modules could act as dynamic filters, evaluating intermediate outputs during the generation process. For example, when generating legal documents, modular systems could enforce compliance with statutory requirements by rejecting outputs that violate predefined legal constraints. This dynamic interaction significantly enhances the reliability and trustworthiness of LLMs, making them suitable for applications where accuracy is paramount.
@@
To summarize, the synergy between MML and LLMs leads to emergent properties that neither system can achieve independently. Inspired by cognitive dual-process theories [11], this hybrid system enables intuitive, fast decision-making through LLMs while ensuring deliberate, logical reasoning via MML-based modules. This dual-layered approach significantly improves decision accuracy and adaptability across diverse applications.
In addition, the interplay between LLM-driven content generation and MML-based validation is able to foster a robust iterative process for hypothesis generation and verification.
By structuring LLMs within the MML framework, AI systems can achieve enhanced robustness, interpretability, and adaptability, setting the foundation for the next generation of unified intelligent systems [12].
4 A Feasible Implementation of MML
We present the methodology of MML, and show the general framework with VQA as an example for instantiation in Fig. 2. In this section, we continue to elaborate on the feasible implementation via introducing three complementary strategies: disentangled representation learning (DRL), neural architecture search (NAS), and neuro-symbolic learning (NSL). First, DRL isolates independent semantic factors into modular representations, promoting transparency, generalizability, and controllability. The disentanglement enables the separation of complex data attributes, making the learned representations more interpretable and robust. Second, NAS automates the design of modular networks by efficiently exploring large search spaces, and identifying optimal operator configurations and interconnections that strike a balance between performance and computational efficiency. This automation significantly reduces human effort in model design while enhancing flexibility. Third, NSL integrates symbolic logic into the inference process within LLMs in a modular manner, enabling structured, rule-based validation and iterative refinement of neural outputs. This symbolic design mitigates issues such as hallucinations and reinforces logical consistency, becoming crucial for high-stakes applications. Collectively, these interconnected strategies exemplify the practical implementation of MML. They not only address the inherent limitations of conventional LLMs but also pave the way for the next generation of robust, interpretable, and logic-capable models in various real-world applications. Fig. 3 illustrates the overall optimization details of this MML implementation.
@@ 4.1 Disentangled Representation Learning for Modular Representation
Definition: Disentangled Representation Learning (DRL)
Let represent a set of disentangled representation modules, where each module maps the input data to a latent variable through a function , i.e.,
A disentangled embedding (a.k.a. representation) denotes the latent variables correspond to statistically independent factors of variation in the input data, such that:
This indicates that each latent variable is responsible for capturing only one underlying factor and remains invariant to changes in other factors (). The modular representation function can then be expressed as:
This function outputs a set of disentangled and semantically meaningful features.
The objective of DRL is to learn disentangled representations by minimizing the following loss function:
@@
This modular design highlights how DRL decomposes complex representations into independent, meaningful components, facilitating robust and interpretable learning. Specifically, DRL enables machine learning models to identify and disentangle hidden factors within observed data, thereby aligning learned representations with real-world semantics. These representations are invariant to external semantic changes [15, 16], aligned with real semantics, and robust against confounding or biased information [17], making them suitable for diverse downstream tasks.
Furthermore, DRL contributes to enhanced interpretability and efficiency in foundation models, such as ChatGPT and Stable Diffusion, by disentangling task-relevant knowledge from redundant components [18]. This capability aids in making foundation models more transparent and adaptable, addressing challenges related to task specificity and interpretability.
@@ 4.2 Neural Architecture Search for Modular Model
where is the -th block of the -th layer, is the weight of the -th block of the -th layer, and is the input to the -th layer, is the number of blocks in the each layer. We can adopt a stack of layers to compose a module,
where is the total number of layers.
Let the architecture search space consist of all possible configurations formed by combining these blocks. Let denote a specific architecture to form a module to achieve a specific function.
where determines how the module weighs the blocks in each layer, and the group of them, i.e. , is the configuration of the module that determines how it selects different calculation paths within the supernetwork to provide the flexibility of the model design.
@@
where is the chosen architecture and is the optimal learned architecture, denotes the loss function that measures the discrepancy between the predicted and the true outputs. The search process optimizes both the modules’ design and their interconnections using search algorithms such as reinforcement learning, evolutionary algorithms, or gradient-based optimization.
This modular design allows for the flexible combination of different neural components while efficiently identifying the optimal architecture, balancing between performance and computational cost. The NAS process typically comprises three key components: search space, search strategy, and evaluation strategy [19].
@@ 4.3 Neuro-Symbolic Learning for Modular Reasoning
Let represent a set of symbolic reasoning modules, where each symbolic module processes a symbolic representation to produce a logical conclusion through a reasoning function , i.e.,
The objective of NSL is to learn a joint modular model that combines neural modules and symbolic modules as follows:
Here, the symbol denotes the modular aggregation operation, which combines the outputs from multiple modules into a unified representation. Specifically, the inner aggregation collects the outputs from the neural modules, while the outer aggregation combines the logical conclusions from the symbolic reasoning modules.
The model’s optimization objective is to minimize the loss function:
@@
This modular design fosters a balanced interaction between neural and symbolic systems. For instance, the “System 1 and System 2” framework by Yoshua Bengio 111https://yoshuabengio.org/2023/03/21/scaling-in-the-service-of-reasoning-model-based-ml/ outlines a collaborative model, where neural modules perform fast pattern recognition while symbolic modules conduct logical reasoning [31]. The modular reasoning is able to optimize both neural networks and symbolic systems, enhancing the efficiency and interpretability of complex reasoning tasks [34]. NSL also shows potential in improving LLMs by integrating logical consistency, guiding inference strategies, and enhancing prompt engineering, making LLMs more robust and interpretable [35, 36].
@@ 5 Challenges and Future Directions
Evaluation and Benchmarking. Traditional evaluation metrics for LLMs—centered on language fluency and task-specific accuracy—fall short of capturing the logical consistency and interpretability gains afforded by MML. Without standardized benchmarks that assess both performance and reasoning quality, it is difficult to measure progress across different systems. Establishing comprehensive evaluation protocols that include metrics for logical coherence, safety, fairness, and explanation quality will be essential. Such benchmarks will not only facilitate rigorous comparisons but also drive the development of models that are robust in safety-critical and high-stakes applications.
@@
In summary, addressing these challenges requires a holistic research agenda that improves both the integration mechanisms and the training methodologies of MML-based LLMs. By developing adaptive interfaces, novel joint optimization techniques, scalable architectures, and standardized evaluation benchmarks, as well as finding effective ways to balance interpretability with performance, the community can unlock the full potential of MML systems. Such advancements will be critical in paving the way toward more robust, transparent, and versatile foundation models that can better meet the demands of real-world applications.
@@ 6 Conclusions
In this paper, we propose a new concept of Modular Machine Learning (MML), elaborating the pivotal role of MML in advancing LLMs. By dissecting the challenges inherent to current LLMs which range from limitations in logical reasoning and factual consistency to issues of interpretability and scalability, we illustrate how a modular learning paradigm can bridge the gap between raw (continuous) neural perception and structured (discrete) logic reasoning. We show that the integration of MML and LLMs not only mitigates issues such as hallucinations and bias, but also paves the way for more explainable and accountable AI applications in critical domains.
Looking ahead, addressing the challenges of configurable interface, joint optimization, and computational efficiency is crucial. Future research should prioritize the development of adaptive interfaces and advanced hybrid training methodologies, as well as the establishment of standardized benchmarks for rigorously evaluating both linguistic and logical competencies. By continuing to refine these approaches, the AI community can harness the full potential of MML, steering large foundation models toward greater reliability and real-world utility. As the final conclusion, we believe that MML ultimately stands as an indispensable strategy for the evolution of new-generation LLMs.

### v2

@@ 5 The Key Significance of MML for LLMs
Enable Counterfactual Reasoning. One of the most profound limitations of current LLMs is their inability to effectively reason about counterfactuals [120]. Counterfactual reasoning explores hypothetical “what if” scenarios, such as estimating outcomes had a patient received a different treatment. While LLMs excel at pattern-based prediction, they remain ill-suited for tasks demanding logical or causal consistency. MML addresses this gap by integrating modular reasoning systems grounded in rules or causal graphs, thereby enabling coherent counterfactual hypotheses [121]. By decoupling perception from reasoning, MML mitigates statistical biases of neural networks and enforces causal constraints, offering a robust foundation for counterfactual reasoning within LLMs [122].
@@
Eliminate Hallucinations. Hallucinations, factually incorrect or inconsistent outputs, arise from LLMs’ probabilistic training and pose severe risks in domains such as healthcare and law. MML mitigates this by embedding modular reasoning components that enforce factuality through knowledge bases, ontologies, or rules. For instance, a medical module can cross-check LLM recommendations against clinical guidelines. MML also supports iterative refinement, where outputs (e.g., scientific summaries) are validated via domain-specific rules, ensuring logical consistency. Beyond post-processing, modular filters can dynamically monitor and correct outputs during generation, such as enforcing statutory compliance in legal drafting. These mechanisms collectively enhance reliability and make LLMs more trustworthy for high-stakes applications.
@@
The interaction between MML and LLMs leads to emergent properties that neither system can achieve independently. Inspired by cognitive dual-process theories [128], this hybrid system enables intuitive, fast decision-making through LLMs while ensuring deliberate, logical reasoning via MML-based modules. This dual-layered approach significantly improves decision accuracy and adaptability across diverse applications. In addition, the interplay between LLM-driven content generation and MML-based validation is able to foster a robust iterative process for hypothesis generation and verification. By structuring LLMs within the MML framework, AI systems can achieve enhanced robustness, interpretability, and adaptability, setting the foundation for the next generation of unified intelligent systems [129].
6 A Unified Framework of MML for LLM
Fig. 3 presents the general methodology of MML in detail and shows the unified framework with VQA as an example for instantiation. In this section, we continue to elaborate on the feasible implementation via introducing three complementary strategies: disentangled representation learning (DRL), neural architecture search (NAS), and neuro-symbolic learning (NSL). First, DRL isolates independent semantic factors into modular representations, promoting transparency, generalizability, and controllability. The disentanglement enables the separation of complex data attributes, making the learned representations more interpretable and robust. Second, NAS automates the design of modular networks by efficiently exploring large search spaces, and identifying optimal operator configurations and interconnections that strike a balance between performance and computational efficiency. This automation significantly reduces human effort in model design while enhancing flexibility. Third, NSL integrates symbolic logic into the inference process within LLMs in a modular manner, enabling structured, rule-based validation and iterative refinement of neural outputs. This symbolic design mitigates issues such as hallucinations and reinforces logical consistency, becoming crucial for high-stakes applications. Collectively, these interconnected strategies exemplify the practical implementation of MML. They not only address the inherent limitations of conventional LLMs but also pave the way for the next generation of robust, interpretable, and logic-capable models in various real-world applications. Fig. 4 illustrates the overall optimization details of this MML implementation.
@@ 6.1 Disentangled Representation Learning for Modular Representation
Definition 1. (Disentangled Representation Learning (DRL)) Let represent a set of disentangled representation modules, where each module maps the input data to a latent variable through a function , i.e.,
A disentangled embedding (a.k.a. representation) denotes the latent variables correspond to statistically independent factors of variation in the input data, such that: This indicates that each latent variable is responsible for capturing only one underlying factor and remains invariant to changes in other factors (). The modular representation function can then be expressed as:
This function outputs a set of disentangled and semantically meaningful features. The objective of DRL is to learn disentangled representations by minimizing the following loss function:
@@
This modular design highlights how DRL decomposes complex representations into independent, meaningful components, facilitating robust and interpretable learning. Specifically, DRL enables machine learning models to identify and disentangle hidden factors within observed data, thereby aligning learned representations with real-world semantics. These representations are invariant to external semantic changes [24, 131], aligned with real semantics, and robust against confounding or biased information [132], making them suitable for diverse downstream tasks. DRL contributes to enhanced interpretability and efficiency in foundation models, such as ChatGPT and Stable Diffusion, by disentangling task-relevant knowledge from redundant components [133]. This capability aids in making foundation models more transparent and adaptable, addressing challenges related to task specificity and interpretability.
@@ 6.2 Neural Architecture Search for Modular Model
where is the -th block of the -th layer, is the weight of the -th block of the -th layer, and is the input to the -th layer, is the number of blocks in the each layer. We can adopt a stack of layers to compose a module, where is the total number of layers.
Let the architecture search space consist of all possible configurations formed by combining these blocks. Let denote a specific architecture to form a module to achieve a specific function: where determines how the module weighs the blocks in each layer, and the group of them, i.e. , is the configuration of the module that determines how it selects different calculation paths within the supernetwork to provide the flexibility of the model design.
@@
where is the chosen architecture and is the optimal learned architecture, denotes the loss function that measures the discrepancy between the predicted and the true outputs. The search process optimizes both the modules’ design and their interconnections using search algorithms such as reinforcement learning, evolutionary algorithms, or gradient-based optimization. This modular design allows for the flexible combination of different neural components while efficiently identifying the optimal architecture, balancing between performance and computational cost. The NAS process typically comprises three key components: search space, search strategy, and evaluation strategy [12].
@@ 6.3 Neuro-Symbolic Learning for Modular Reasoning
Let represent a set of symbolic reasoning modules, where each symbolic module processes a symbolic representation to produce a logical conclusion through a reasoning function , i.e., The objective of NSL is to learn a joint modular model that combines neural modules and symbolic modules as follows:
Here, the symbol denotes the modular aggregation operation, which combines the outputs from multiple modules into a unified representation. Specifically, the inner aggregation collects the outputs from the neural modules, while the outer aggregation combines the logical conclusions from the symbolic reasoning modules. The model’s optimization objective is to minimize the loss function:
@@
This modular design fosters a balanced interaction between neural and symbolic systems. For instance, the “System 1 and System 2” framework by Yoshua Bengio 111https://yoshuabengio.org/2023/03/21/scaling-in-the-service-of-reasoning-model-based-ml/ outlines a collaborative model, where neural modules perform fast pattern recognition while symbolic modules conduct logical reasoning [142]. The modular reasoning is able to optimize both neural networks and symbolic systems, enhancing the efficiency and interpretability of complex reasoning tasks [7]. NSL also shows potential in improving LLMs by integrating logical consistency, guiding inference strategies, and enhancing prompt engineering, making LLMs more robust and interpretable [145].
@@ 7 Challenges and Future Directions
Evaluation and Benchmarking. Traditional evaluation metrics for LLMs, centered on language fluency and task-specific accuracy, fall short of capturing the logical consistency and interpretability gains afforded by MML. Without standardized benchmarks that assess both performance and reasoning quality, it is difficult to measure progress across different systems. Establishing comprehensive evaluation protocols that include metrics for logical coherence, safety, fairness, and explanation quality will be essential. Such benchmarks will not only facilitate rigorous comparisons but also drive the development of models that are robust in safety-critical and high-stakes applications.
@@
In summary, advancing MML-based LLMs requires improving integration mechanisms and training methodologies through adaptive interfaces, efficient joint optimization, scalable architectures, and standardized benchmarks, while balancing interpretability with performance. Such progress will be essential for building more robust, transparent, and versatile foundation models suited to real-world applications.
@@ 8 Conclusions
The rapid development of LLMs has revealed both their transformative potential and their limitations in reasoning, interpretability, and adaptability. This survey provides a systematic review of Modular Machine Learning (MML), organizing along two major dimensions: Modular Data Representation and Modular Model Optimization. We examine representative approaches under supervised and unsupervised settings, ranging from representation disentanglement to modular networks, symbolic constraints, neural architecture search, and inductive modularization, highlighting how modular principles enhance transparency, reliability, scalability, and extensibility compared to monolithic systems. Building on this taxonomy, we propose a unified framework of modular representation, modular model, and modular reasoning, as well as discuss its significance for advancing next-generation LLMs. Looking ahead, open challenges include designing configurable interfaces, achieving efficient joint optimization, improving scalability, and establishing standardized benchmarks. In conclusion, we would like to argue that MML offers both a unifying perspective and a promising paradigm for guiding the evolution of reliable, interpretable, and adaptable large language models.